An object starts from rest with a constant acceleration of 10 ft/s² along a straight road. Determine the velocity after 3.5 sec, b) the average velocity for the 3.5 seconds interval, and c) the distance travelled in this given time.


1
Expert's answer
2022-03-18T09:08:24-0400

(a)

"v=v_0+at\\
=\rm 0\: ft/s+10\: ft/s^2\times 3.5\: s=35\: ft/s"

(b)

"v_{\rm ave}=\frac{v_0+v}{2}\\
=\rm\frac{0\: ft/s+35\: ft/s}{2}=17.5\: ft/s"

(c)

"d=v_{\rm ave}t\\
=\rm 17.5\: ft/s\times 3.5\: s=61.25\: ft"